OpenStudy (pepe_the_frog):

48 ÷ ((2 + 6) · 2) - 1

alones (alones):

Alright I'm nto sure maybe this how it would look \[\frac{ 48 }{ ((2 + 6) · 2) - 1 }\]

OpenStudy (pepe_the_frog):

i see.. hm one sec

alones (alones):

\[\frac{ 48 }{ (8)(2)}-1\] \[\frac{ 48 }{ 16 }-1\] \[\ 3−1\] I got \( 2\) as my answer.

OpenStudy (shansha1):

4816−1 3−1 so ur answer is 2

OpenStudy (shansha1):

48 divide by 16

OpenStudy (mathmale):

This problem is a test of your knowledge of order of operations rules (PEMDAS). In this particular problem you must multiply and divide BEFORE you add or subtract. Give the problem another try, please, showing your work step by step.